OpenStudy (alayrhi):

The price of a fan is $15. The price is 27% lower than last week. What was the price of the fan last week? Round your answer to the nearest cent.

OpenStudy (mathstudent55):

The price last week was 100% of that price. The price this week is 27% lower than the price last week. What is 100% minus 27%?

OpenStudy (mathstudent55):

100 - 27 = ?

OpenStudy (alayrhi):

73 @mathstudent55

OpenStudy (mathstudent55):

Good. 100 - 27 = 73, so 100% - 27% = 73% That means that the $15 you pay this week is 73% of the original price.

OpenStudy (mathstudent55):

The original price is unknown, so we call it x. The price this week is 73% of the original price, and it is $15. We can think of this problem as the following problem: 73% of what number is 15 ? We write an equation based on this question. 73% of x = 15 In math, "of" usually means ""times": 73% * x = 15 Also, % means "divided by 100." \(\dfrac{73}{100} x = 15\) or \(0.73x = 100\) Can you solve this equation (or the one above with the fraction) for x?

OpenStudy (alayrhi):

Um x equals 136.986301369863

OpenStudy (alayrhi):

I think

OpenStudy (mathmale):

Does $136 sound reasonable, if the original price were somewhere around $15?

OpenStudy (alayrhi):

Nah

OpenStudy (mathmale):

Let x represent the price of the fan last week, before the sale started. This week's price is x - 0.27x = $15. Solve this for x, please.

OpenStudy (alayrhi):

I guessing the answer is no higher than 50

OpenStudy (mathmale):

Guessing has no place here, sorry.

OpenStudy (alayrhi):

:|

OpenStudy (alayrhi):

Well okay

OpenStudy (mathstudent55):

$136 is too much. You have an an equation in which a number times x equals another number. You want x alone, so just divide both sides by the number multiplying x. \(0.73x = 15\) \(\dfrac{0.73}{\color{red}{0.73}} x = \dfrac{15}{\color{red}{0.73}} \) What is 15/0.73?